OpenStudy (anonymous):

-16x^2+9y^2+32x+144y-16=0 Find the standard formula

OpenStudy (anonymous):

@Directrix

OpenStudy (agent0smith):

\[\large -16x^2+9y^2+32x+144y-16=0 \] First step is to group all the common terms together, and maybe add 16 to both sides... \[\large (-16x^2+32x)+(9y^2+144y)= 16\] Now you try completing the square, first you need to factor out the -16 from the x^2 and the 9 from the y^2: \[\large -16(x^2-2x)+9(y^2+16y)=16\]